# Towards Agentic Intelligence for Materials Science

### v1

@@ 2.2 Transformer and Foundation Models
To address the computational efficiency bottleneck of Transformer-backed architectures on long scientific sequences, several alternative architectures have been proposed. Structured State Space models (S4) 140 addressed the limitations of time-series materials data and molecular dynamics simulations by implementing parameter-efficient architectures that facilitated training on exceptionally lengthy sequences (16K+ tokens). Mamba 139, a selective state space model with input-dependent state transitions and linear computing complexity, can reach Transformer-level quality on language-modeling benchmarks while scaling more favorably with sequence length. This efficiency is particularly relevant when models must handle long molecule strings(e.g., SMILES or SELFIES for complex molecules and polymers), simulation trajectories from molecular dynamics or DFT relaxations, lab notebooks accumulated over long experimental campaigns, or tool outputs as part of an end-to-end discovery workflow. The next paradigm shift moves away from passively reading text and toward reasoning and tool usage skills, leading to autonomous agents.
2.3 Large Language Models and Agents
The recent innovations in AI have introduced what 324 is termed “The Era of Experience”. In this view, progress is driven not only by fitting static human-generated corpora, but also by iterative interaction with environments that provide feedback signals. Now the systems-based LLM can complete complex experimental tasks to be an agent rather than a single network predictor. By interacting with the work, the agent can learn, improve, and adapt to real-world tasks.
The AI field is currently transitioning from the “Era of Human Data,” where progress relies on imitating finite, static datasets, to the “Era of Experience,” where agents achieve superhuman capabilities by actively interacting with their environments and learning from the consequences of their actions 324. The relevance of this shift to materials science is practical: the interactive nature can steer the whole pipeline towards the ultimate goal of discovering novel and useful mateirals.
Just as general-purpose agents are now moving beyond human-privileged text interactions to autonomous streams of experience, AI for materials science must evolve from passive models trained on historical property databases to active Scientist AI 372 agents that engage in closed-loop discovery. By treating experiments and simulations not merely as data sources but as interactive environments, such agents can generate their own high-quality data through reinforcement learning, potentially expanding exploration beyond human-selected candidates and accelerating iterative design–test–learn cycles.
@@ 2.3 Large Language Models and Agents
New paradigms, such as alignment and instruction tuning (e.g., Constitutional AI 29, RLHF 267), have improved instruction following and controllability and can reduce hallucinations in practice, although factuality is not guaranteed and typically depends on grounding and verification, especially for scientific domains like materials science. These paradigms are based on foundation models and can be viewed as post-training objective shaping on top of pretraining. Few-shot 112 and zero-shot learning 292 have made it easier to adapt tasks with little data. This is important in fields where labeled datasets are limited by the high cost of experiments and the need for domain expertise. In-context learning 56, 121 enables rapid prototyping of new actions directly from stimuli, removing the need for task-specific fine-tuning. In order to improve scientific decision-making rather than just isolated perception, models are becoming more skilled at tool- and evidence-grounded reasoning across a variety of modalities, such as images 72, crystal structures 393, spectroscopic data, and experimental measurements. These works suggest that agentic systems couple language-based reasoning with external actions (e.g., retrieval, simulation, and code execution).
@@
Enhanced agentic systems have been developed during the evolution of LLMs. In order to make decisions and improve grounding, agent loops combine reasoning and action by utilizing external tools such as retrieval 208, calculators and symbolic computation 119, code execution or function-calling 403, browsers 254, structured databases or enterprise APIs 290. Later frameworks like Toolformer 309, AutoGPT 135, and LangChain 67 used these features to add persistent memory, the ability to work with more than one tool at a time, and the ability to plan recursively. These works demonstrate that the agentic system can address problems necessitating extended procedures, although consistent long-term performance continues to be influenced by tool reliability, feedback design, and safety constraints.
@@
In the second half of 2025, a plethora of agents with thinking and embodied action have emerged. Anthropic’s Claude Sonnet 24 is a cognitive agent equipped with enhanced reasoning, understanding, and tool-using abilities. This is a step toward AI systems that are better at using tools and sustaining multi-step analysis in real workflows. It also supports more scientific use by helping users structure questions, integrate evidence, and connect ideas across domains. Gemini Robotics 1.5 122 demonstrated transfer learning across different robotic embodiments, advancing robotic agents beyond single-platform policies. The vision-language-action (VLA) model shows that it is smart by making its thoughts clear before it does anything. NVIDIA’s Cosmos reasoning model 260 improves agent reasoning by integrating physical priors and commonsense constraints into robotic decision-making. Microsoft’s autonomous agents317 demonstrate that AI agents have great potential to orchestrate complex, multi-step tasks on behalf of users.
Through the above improvements, multiple agents with different skills can collaborate to build agent ecosystems capable of completing complex tasks. Recent systems demonstrate that AI can function as a powerful research partner, able to read literature, generate hypotheses, plan experiments, and design syntheses, rather than just making predictions 231. Such agents can retrieve external resources, call specialized tools, and coordinate downstream actions autonomously 228, 135. Integrating data-driven analysis, such as extracting useful information from huge amounts of scientific literature, with formulating hypotheses and organizing experiments can be achieved directly by LLMs. Such capabilities are discussed in detail in 164. These features make it much easier to sort data for tasks that only apply to a certain area. They also build independent, agent-driven research pipelines that speed up the process of making new discoveries by emphasizing system-level integration (feedback, coordination, and objective alignment) rather than isolated model capabilities.
@@ 2.4 Pipeline-Centric Perspective
After addressing Q1, we discuss the interaction between the pre‑trained LLMs and the agentic LLMs, as shown in Fig. 4, from the end‑to‑end pipeline‑centric perspective, explicitly treating the pretraining stage as a controllable module whose data, objectives, and safety constraints can be continuously revised in light of downstream discovery performance rather than as a one‑off static initialization. This framing makes pretraining an object of system design—subject to iterative revision—rather than a fixed prerequisite that is optimized once and then taken for granted.
@@
It is generally believed that general pre‑training corpora on which contemporary LLMs are trained are essential to the downstream fine‑tuned models; however, recent studies (e.g., 172) suggest that such broad pretraining may not always be necessary and, in some cases, may even be suboptimal for specialized scientific applications. Furthermore, reliance on indiscriminately scraped web‑scale corpora increases the chance of poisoning attacks and incidental misinformation, which can induce harmful or misleading behaviors even when benchmarks appear unaffected 331. Instead, if the ultimate objective of discovering novel materials is what we value most, we should review the techniques presented before and enforce that every upstream design choice—including corpus selection, pretraining objectives, and alignment procedures—is justified by its contribution to this end goal rather than by improvements on proxy language benchmarks. More concretely, better pretraining should be defined by downstream discovery-relevant behaviors (e.g., planning feasible syntheses and correctly handling hazards), not by generic language metrics alone. Continuing the presentations in Sec. 2.2, we should require that these architectures benefit the discovery of novel materials and not just improve the pre‑training scores on generic benchmarks, recognizing that even small fractions of poisoned or misaligned data can strongly bias downstream scientific behavior.
@@
To make this pipeline‑centric stance operational, the pretraining stage must be embedded within a feedback loop in which agentic systems operating over simulations and experiments supply long‑horizon reward signals that can, in principle, be propagated back to earlier stages to revise corpus composition, sampling strategies, and alignment objectives whenever inherited behaviors of a general‑purpose LLM systematically hinder materials discovery. This suggests moving from a frozen pretraining corpus toward an adaptive, feedback‑driven pretraining loop where materials‑science outcomes and safety considerations jointly determine which segments of the original web‑scale data are amplified, attenuated, or removed, and where domain‑specific corpora are dynamically expanded in response to identified capability gaps along the design–synthesis–validation pipeline. Recent progress in LLM‑based agent architectures and tooling, as discussed in Sec. 2.3, makes such a pipeline‑centric perspective increasingly feasible, allowing the successful discovery of novel, experimentally validated materials to serve as the primary reward signal guiding the entire system rather than surrogate benchmarks alone. Crucially, this closes the evaluation loop: it ties what we optimize (training choices) to what we ultimately value (validated discovery).
@@
Influence functions 218, 136 offer one concrete mechanism for such backward credit assignment by estimating how infinitesimal upweighting or downweighting of individual training examples would perturb a model’s parameters and downstream predictions. Approximate influence estimators 218 based on scalable inverse‑Hessian approximations have been demonstrated on large LLMs, recovering sparse, interpretable sets of pretraining sequences that are most responsible for particular emergent behaviors. From a pipeline‑centric perspective, these tools can be repurposed to identify which general‑domain documents and which MatSci‑specific corpora most strongly support or undermine key capabilities needed for discovery (e.g., robust reasoning over phase diagrams, synthesis planning under realistic process constraints, or safety‑critical hazard identification), thereby informing which data sources to emphasize, de‑emphasize, or excise in subsequent rounds of pretraining. This turns data selection into a goal-driven hyperparameter, rather than a one-time curation decision.
@@
Analogously, an AI4MatSci pipeline could employ influence scores computed on discovery‑relevant behaviors, e.g., successful identification of synthesizable candidates, safe handling of hazardous chemistries, or robust performance in long‑horizon design loops. It can upweight beneficial segments of both general and MatSci‑specific data, while downweighting or removing sequences that systematically induce hallucinated mechanisms or impossible synthesis conditions. In this view, pretraining, domain adaptation, and instruction tuning collectively form a single, influence‑addressable memory that is continuously reshaped by reward signals originating from agentic closed loops in simulation and experiments, rather than being optimized only once against proxy language benchmarks that may fail to reveal poisoning or subtle misalignment. In this sense, influence functions become not only interpretability tools but also prospective levers for dynamically re‑aligning the pretraining component with the ultimate materials‑science objective of discovering novel, useful, and safe materials, while mitigating security risks introduced by open‑world data collection. At the same time, these general‑purpose agents must be carefully tailored to materials science to align with real workflows, safety regulations, and experimental constraints, a refinement that will be developed in the next section.
@@ 3 Reactive Tasks in Materials Science from an AI Perspective
298 have emphasized the importance of leveraging AI to accelerate the discovery of advanced materials technologies by integrating information and decision-making across the composition–process–structure–property–performance chain. Having discussed recent AI progress in the previous section, we will now review the evolution of traditional AI for materials science here to ultimately answer the second guiding question Q2 introduced in Sec. 1: Why are existing AI for materials science systems not enough to meet the needs of autonomous materials discovery, even though they score well on material-related benchmarks?
@@ 3.1 Prediction
Applying LLMs to materials science marks a significant paradigm shift, moving predictive modeling from structure-specific deep learning toward generalizable, knowledge-intensive representation learning 207. However, most current deployments still operate in a largely reactive, task-conditioned manner. Prediction in this context encompasses both quantitative forecasting of continuous properties (regression) and categorical assignment (classification), applied across diverse materials data modalities. We argue that the key innovation of applying LLMs for materials science lies in its dual capacity to process symbolic structural inputs (e.g., formulas 425, space-group/Wyckoff descriptors 181, and other text-encoded crystallographic representations 25) and high-dimensional feature sets (e.g., Crystal–structure–derived geometric features 330, DFT-derived features 115, foundation-model embeddings 345). This capability not only helps the development of predictive models with enhanced performance and multimodal adaptability but also facilitates the assessment of their trustworthiness 248.
@@ 3.1.1 Regression Tasks
Regression tasks for materials predict continuous-valued physical properties, such as thermodynamic stability 311, mechanical stiffness 150, and electrical and thermal characteristics 416. This is a crucial step in accelerating the search for and design of novel materials. In practice, these tasks typically assume fixed property definitions and evaluation protocols, which can limit adaptability in end-to-end discovery pipelines 301. The effectiveness of these prediction models relies on their contextually enriched embeddings, which are obtained from various representations of the materials, such as crystal structure 393, chemical composition 130, or textual descriptions 358. Standard machine learning and GNN models have established strong benchmarks. However, the deployment of LLMs has introduced improved capabilities for property prediction by utilizing global context and multi-modal data 207. Still, it has trouble adapting 12 across domain shift, new target properties, and constraints relevant to synthesis and processing.
The performance of machine learning models critically depends on their input representations. This challenge was acutely evident in early models for periodic solids, which had difficulty formulating representations that were invariant to arbitrary choices of coordinate systems or unit cells while guaranteeing uniqueness 314. Early approaches adapted representations from molecular machine learning (e.g., the Coulomb matrix) to periodic systems by incorporating Ewald sums or considering neighboring unit cells 106. A paradigm shift was subsequently established by GNNs through representing crystals as graphs with atoms and bonds as nodes and edges, which allows models to learn from the local atomic environment through message-passing 394. More recently, Transformer-based architectures have built upon this graph-based foundation by processing materials as sequences of tokens—whether atoms, sites, or textual descriptions—a strategy that, by focusing on global context, more effectively captures long-range interactions and symmetries than the locally oriented GNNs leading to improved modeling performance 333. However, improved representations alone do not resolve long-horizon planning, feedback coupling, or decision-level constraints that arise when predictors are embedded in closed-loop discovery workflows.
@@
In current benchmark-driven screening workflows, much emphasis has been placed on the prediction of fundamental electronic properties, especially the bandgap and conductivity for the screening of semiconductors and functional materials 100. This focus is partly pragmatic: these targets are widely used as first-pass filters, but they can also over-represent what is easiest to score on proxy benchmarks. Recent works based on Transformers have shown a superior capability in capturing global crystal characteristics compared to previous successes of GNNs at capturing localized atomic environments 170. The CrystalTransformer model embodies this advance by feeding in crystal structure information through a Transformer encoder; specifically, this work starts with a concatenation of chemical and structural feature matrices, and then processes the combined representation into a Transformer encoder. In the encoder, the multi-head self-attention identifies complex, long-range correlations between all atoms in a manner that is agnostic to their spatial proximity, enabling the model to learn highly context-aware atomic embeddings 169. Other variants, such as MatInFormer, have generalized this paradigm by learning the ”grammar of crystallography” through the tokenization of space group information and subsequently demonstrating the adaptability of this architecture for property prediction on a range of datasets 217. Increasingly, a key factor is the strategic use of linguistic input. For example, if natural language descriptions of a crystalline solid are used, then high-accuracy predictions of properties such as bandgap can be achieved, which seems not possible for some GNNs due to an inability to capture certain nuanced, global structural details such as space group symmetry and Wyckoff positions 304. The success of Transformer-based architectures in the processing of both symbolic structural information and text suggests that they operate as superior contextual encoders rather than purely topological encoders 350. This allows them to implicitly integrate learned chemical and structural knowledge, avoiding the limitations of localized message-passing schemes.
@@
The prediction of mechanical properties such as bulk modulus and shear modulus is important in material design with targeted strength, durability, and resilience 171, 178. Various machine learning models have been widely adopted for predicting such mechanical properties, greatly reducing their dependence on simulation methods that are computationally intensive 206. More recently, domain-specific LLMs for materials science can be specialized in the predictions of complex mechanical properties, often by leveraging text-derived representations or hybrid text+structured inputs. For example, ElaTBot is an LLM for the prediction of the full elastic constant tensor from text inputs 222. Hence, the elastic tensor can be predicted simultaneously with the determination of properties such as bulk modulus at any finite temperature. Moreover, by incorporating general-purpose LLMs like GPT-4o and using the Retrieval-Augmented Generation technique, its prediction accuracy can be boosted, indicating again the power of combining domain-specific knowledge with the broad reasoning capabilities of generalist models when retrieval sources are curated and verifiable 224. This significantly accelerates the prediction process but, more importantly, can support early-stage inverse design by proposing candidate compositions with targeted elastic properties—subject to the fidelity of the predictive surrogate and downstream verification.
@@
Among the most basic and important thermodynamic properties is the formation energy, which dictates stability for a crystalline solid 62. Its prediction is thus of paramount importance in order for high-throughput screening to be performed on new, synthesizable materials. Machine learning models have thus been employed to directly predict it from crystal structures, with the key ingredient being sophisticated feature engineering to account for periodicity in solids 106. Subsequent deep learning models demonstrated the feasibility of material property inference from merely elemental composition and symmetry classification 151.
A particularly noticeable development in this domain has been the development of structure-agnostic models, which predict properties from stoichiometry alone. Using a message-passing architecture on a dense weighted graph representation of the chemical formula, the Roost framework learns material descriptors in an explicit-structure-free way 131. This is especially important in the early stage of materials discovery when little to no crystal structural information is known. Self-supervised pre-training methods have been shown to boost the performances of such models even further, with substantial improvements in the low-data regime 90. Furthermore, LLMs have been demonstrated to handle simple chemical formulas and predict complex chemical properties (e.g., formation energy), significantly reducing the computational cost of material screening.
@@
Thermoelectric materials enable the conversion between thermal and electrical energy, making them crucial for waste heat recovery and solid-state cooling technologies. However, the lack of large, structured datasets with relevant performance metrics hinders the development of models for high-performance thermoelectric materials 377. LLM-based extraction approaches are increasingly used to mitigate this data bottleneck. Recent work has shown that agentic, LLM-driven workflows autonomously extract the thermoelectric and structural properties from thousands of scientific articles 126. These intelligent agents can parse text, tables, and captions to create large machine-readable datasets that couple performance metrics with structural information such as the crystal class and space group 192. This automated curation pipeline can improve downstream regression by expanding training data and improving schema consistency; however, it still primarily optimizes proxy prediction objectives unless coupled to closed-loop validation and decision-making in the discovery loop.

### v2

@@ 2.2 Transformer and Foundation Models
To address the computational efficiency bottleneck of Transformer-based architectures on long scientific sequences, several alternative architectures have been proposed. Structured State Space Models (S4) 141 addressed the limitations of time-series materials data and molecular dynamics simulations by implementing parameter-efficient architectures that facilitated training on exceptionally lengthy sequences (16K+ tokens). Mamba 140, a selective state space model with input-dependent state transitions and linear computing complexity, can reach Transformer-level quality on language-modeling benchmarks while scaling more favorably with sequence length. This efficiency is particularly relevant when models must handle long molecular strings (e.g., SMILES or SELFIES for complex molecules and polymers), simulation trajectories from molecular dynamics or density functional theory (DFT) relaxations, lab notebooks accumulated over long experimental campaigns, or tool outputs as part of an end-to-end discovery workflow. The next paradigm shift moves away from passively reading text and toward reasoning and tool usage skills, leading to autonomous agents.
2.3 LLMs and Agents
The recent innovations in AI have introduced what is termed “The Era of Experience” 326. In this view, progress is driven not only by fitting static human-generated corpora, but also by iterative interaction with environments that provide feedback signals. Now the systems-based LLM can complete complex experimental tasks to be an agent rather than a single network predictor. By interacting with the environment/workflow, the agent can learn, improve, and adapt to real-world tasks.
The AI field is currently transitioning from the “Era of Human Data,” where progress relies on imitating finite, static datasets, to the “Era of Experience,” where agents achieve superhuman capabilities by actively interacting with their environments and learning from the consequences of their actions 326. The relevance of this shift to materials science is practical: the interactive nature can steer the whole pipeline towards the ultimate goal of discovering novel and useful materials.
Just as general-purpose agents are now moving beyond human-privileged text interactions to autonomous streams of experience, AI for materials science must evolve from passive models trained on historical property databases to active AI Scientist 375 agents that engage in closed-loop discovery. By treating experiments and simulations not merely as data sources but as interactive environments, such agents can generate their own high-quality data through RL, potentially expanding exploration beyond human-selected candidates and accelerating iterative design–test–learn cycles.
@@ 2.3 LLMs and Agents
New paradigms, such as alignment and instruction tuning (e.g., Constitutional AI 30, RLHF 269), have improved instruction following and controllability and can reduce hallucinations in practice, although factuality is not guaranteed and typically depends on grounding and verification, especially for scientific domains like materials science. These paradigms are based on foundation models and can be viewed as post-training objective shaping on top of pre-training. Few-shot 113 and zero-shot learning 294 have made it easier to adapt to tasks with little data. This is important in fields where labeled datasets are limited by the high cost of experiments and the need for domain expertise. In-context learning 57, 122 enables rapid prototyping of new actions directly from stimuli, removing the need for task-specific fine-tuning. In order to improve scientific decision-making rather than just isolated perception, models are becoming more skilled at tool- and evidence-grounded reasoning across a variety of modalities, such as images 73, crystal structures 396, spectroscopic data, and experimental measurements. These works suggest that agentic systems couple language-based reasoning with external actions (e.g., retrieval, simulation, and code execution).
@@
Enhanced agentic systems have been developed during the evolution of LLMs. In order to make decisions and improve grounding, agent loops combine reasoning and action by utilizing external tools such as retrieval 209, calculators and symbolic computation 120, code execution or function calling 406, browsers 256, structured databases or enterprise APIs 292. Later frameworks like Toolformer 311, AutoGPT 136, and LangChain 68 used these features to add persistent memory, the ability to work with more than one tool at a time, and the ability to plan recursively. These works demonstrate that agentic systems can address problems necessitating extended procedures, although consistent long-term performance continues to be influenced by tool reliability, feedback design, and safety constraints.
@@
In the second half of 2025, a wave of agents with thinking and embodied action have emerged. Anthropic’s Claude Sonnet 25 is a cognitive agent equipped with enhanced reasoning, understanding, and tool-using abilities. This is a step toward AI systems that are better at using tools and sustaining multi-step analysis in real workflows. It also supports more scientific use by helping users structure questions, integrate evidence, and connect ideas across domains. Gemini Robotics 1.5 123 demonstrated transfer learning across different robotic embodiments, advancing robotic agents beyond single-platform policies. Vision-language-action (VLA) models highlight the value of explicit intermediate plans or reasoning steps that can improve action. NVIDIA’s Cosmos reasoning model 262 improves agent reasoning by integrating physical priors and common sense constraints into robotic decision-making. Microsoft’s autonomous agents 319 demonstrate that AI agents have great potential to orchestrate complex, multi-step tasks on behalf of users.
Through the above improvements, multiple agents with different skills can collaborate to build agent ecosystems capable of completing complex tasks. Recent systems demonstrate that AI can function as a powerful research partner, able to read literature, generate hypotheses, plan experiments, and design syntheses, rather than just making predictions 233. Such agents can retrieve external resources, call specialized tools, and coordinate downstream actions autonomously 230, 136.
Integrating data-driven analysis, such as extracting useful information from large-scale scientific literature, with formulating hypotheses and organizing experiments can be supported by LLM-based agents. Such capabilities are discussed in detail in 165. These features make it much easier to sort data for tasks that only apply to a certain area.
However, end-to-end materials discovery is still limited by reliable grounding, constraint-aware reasoning such as stability and synthesizability, and closed-loop validation with simulations and experiments. Agents can also enable more independent, agent-driven research pipelines that may accelerate discoveries by emphasizing system-level integration (feedback, coordination, and objective alignment) rather than isolated model capabilities.
@@ 2.4 Pipeline-Centric Perspective
After addressing Q1, we discuss the interaction between the pre‑trained LLMs and the agentic LLMs, as shown in Fig. 4, from the end‑to‑end pipeline‑centric perspective, explicitly treating the pre-training stage as a controllable module whose data, objectives, and safety constraints can be continuously revised in light of downstream discovery performance rather than as a one‑off static initialization. This framing makes pre-training an object of system design—subject to iterative revision—rather than a fixed prerequisite that is optimized once and then taken for granted.
@@
It is generally believed that general pre-training corpora on which contemporary LLMs are trained are essential to the downstream fine-tuned models; however, recent studies (e.g., 173) suggest that such broad pre-training may not always be necessary and, in some cases, may even be suboptimal for specialized scientific applications. Furthermore, reliance on indiscriminately scraped web‑scale corpora increases the chance of poisoning attacks and incidental misinformation, which can induce harmful or misleading behaviors even when benchmarks appear unaffected 333. Instead, if the ultimate objective of discovering novel materials is what we value most, we should review the techniques presented before and enforce that every upstream design choice—including corpus selection, pre-training objectives, and alignment procedures—is justified by its contribution to this end goal rather than by improvements on proxy language benchmarks. More concretely, better pre-training should be defined by downstream discovery-relevant behaviors (e.g., planning feasible syntheses and correctly handling hazards), not by generic language metrics alone. Continuing the presentations in Sec. 2.2, we should require that these architectures benefit the discovery of novel materials and not just improve the pre‑training scores on generic benchmarks, recognizing that even small fractions of poisoned or misaligned data can strongly bias downstream scientific behavior.
@@
To make this pipeline‑centric stance operational, the pre-training stage must be embedded within a feedback loop in which agentic systems operating over simulations and experiments supply long‑horizon reward signals that can, in principle, be propagated back to earlier stages to revise corpus composition, sampling strategies, and alignment objectives whenever inherited behaviors of a general‑purpose LLM systematically hinder materials discovery. This suggests moving from a frozen pre-training corpus toward an adaptive, feedback‑driven pre-training loop where materials‑science outcomes and safety considerations jointly determine which segments of the original web‑scale data are amplified, attenuated, or removed, and where domain‑specific corpora are dynamically expanded in response to identified capability gaps along the design–synthesis–validation pipeline. Recent progress in LLM‑based agent architectures and tooling, as discussed in Sec. 2.3, makes such a pipeline‑centric perspective increasingly feasible, allowing the successful discovery of novel, experimentally validated materials to serve as the primary reward signal guiding the entire system rather than surrogate benchmarks alone. Crucially, this closes the evaluation loop: it ties what we optimize (training choices) to what we ultimately value (validated discovery).
@@
Influence functions 220, 137 offer one concrete mechanism for such backward credit assignment by estimating how infinitesimal upweighting or downweighting of individual training examples would perturb a model’s parameters and downstream predictions. Approximate influence estimators 220 based on scalable inverse‑Hessian approximations have been demonstrated on large LLMs, recovering sparse, interpretable sets of pre-training sequences that are most responsible for particular emergent behaviors. From a pipeline‑centric perspective, these tools can be repurposed to identify which general‑domain documents and which MatSci‑specific corpora most strongly support or undermine key capabilities needed for discovery (e.g., robust reasoning over phase diagrams, synthesis planning under realistic process constraints, or safety‑critical hazard identification), thereby informing which data sources to emphasize, de‑emphasize, or excise in subsequent rounds of pre-training. This turns data selection into a goal-driven hyperparameter, rather than a one-time curation decision.
@@
Analogously, an AI4MatSci pipeline could employ influence scores computed on discovery-relevant behaviors, e.g., successful identification of synthesizable candidates, safe handling of hazardous chemistries, or robust performance in long‑horizon design loops. It can upweight beneficial segments of both general and MatSci‑specific data, while downweighting or removing sequences that systematically induce hallucinated mechanisms or impossible synthesis conditions. In this view, pre-training, domain adaptation, and instruction tuning collectively form a single, influence‑addressable memory that is continuously reshaped by reward signals originating from agentic closed loops in simulation and experiments, rather than being optimized only once against proxy language benchmarks that may fail to reveal poisoning or subtle misalignment. In this sense, influence functions become not only interpretability tools but also prospective levers for dynamically re‑aligning the pre-training component with the ultimate materials‑science objective of discovering novel, useful, and safe materials, while mitigating security risks introduced by open‑world data collection. At the same time, these general‑purpose agents must be carefully tailored to materials science to align with real workflows, safety regulations, and experimental constraints, a refinement that will be developed in the next section.
@@ 3 Reactive Tasks in Materials Science from an AI Perspective
300 have emphasized the importance of leveraging AI to accelerate the discovery of advanced materials technologies by integrating information and decision-making across the composition-process-structure-property-performance chain. Having discussed recent AI progress in the previous section, we will now review the evolution of traditional AI for materials science here to ultimately answer the second guiding question Q2 introduced in Sec. 1: Why are existing AI for materials science systems not enough to meet the needs of autonomous materials discovery, even though they score well on material-related benchmarks?
@@ 3.1 Prediction
Applying LLMs to materials science marks a significant paradigm shift, moving predictive modeling from structure-specific deep learning toward generalizable, knowledge-intensive representation learning 208. However, most current deployments still operate in a largely reactive, task-conditioned manner. Prediction in this context encompasses both quantitative forecasting of continuous properties (regression) and categorical assignment (classification), applied across diverse materials data modalities. We argue that the key innovation of applying LLMs for materials science lies in its dual capacity to process symbolic structural inputs (e.g., formulas 428, space-group/Wyckoff descriptors 182, and other text-encoded crystallographic representations 26) and high-dimensional feature sets (e.g., Crystal–structure–derived geometric features 332, DFT-derived features 116, foundation-model embeddings 347). This capability not only helps the development of predictive models with enhanced performance and multimodal adaptability but also facilitates the assessment of their trustworthiness 250.
@@ 3.1.1 Regression Tasks
Regression tasks for materials predict continuous-valued physical properties, such as thermodynamic stability 313, mechanical stiffness 151, and electrical and thermal characteristics 419. This is a crucial step in accelerating the search for and design of novel materials. In practice, these tasks typically assume fixed property definitions and evaluation protocols, which can limit adaptability in end-to-end discovery pipelines 303. The effectiveness of these prediction models relies on their contextually enriched embeddings, which are obtained from various representations of the materials, such as crystal structure 396, chemical composition 131, or textual descriptions 361. Standard machine learning and GNN models have established strong benchmarks. However, the deployment of LLMs has introduced improved capabilities for property prediction by utilizing global context and multi-modal data 208. Still, it has trouble adapting 12 across domain shift, new target properties, and constraints relevant to synthesis and processing.
The performance of machine learning models critically depends on their input representations. This challenge was acutely evident in early models for periodic solids, which had difficulty formulating representations that were invariant to arbitrary choices of coordinate systems or unit cells while guaranteeing uniqueness 316. Early approaches adapted representations from molecular machine learning (e.g., the Coulomb matrix) to periodic systems by incorporating Ewald sums or considering neighboring unit cells 107. A paradigm shift was subsequently established by GNNs through representing crystals as graphs with atoms and bonds as nodes and edges, which allows models to learn from the local atomic environment through message-passing 397. More recently, Transformer-based architectures have built upon this graph-based foundation by processing materials as sequences of tokens—whether atoms, sites, or textual descriptions—a strategy that, by focusing on global context, more effectively captures long-range interactions and symmetries than the locally oriented GNNs leading to improved modeling performance 335. However, improved representations alone do not resolve long-horizon planning, feedback coupling, or decision-level constraints that arise when predictors are embedded in closed-loop discovery workflows.
@@
In current benchmark-driven screening workflows, much emphasis has been placed on the prediction of fundamental electronic properties, especially the bandgap and conductivity for the screening of semiconductors and functional materials 101. This focus is partly pragmatic: these targets are widely used as first-pass filters, but they can also over-represent what is easiest to score on proxy benchmarks. Recent works based on Transformers have shown a superior capability in capturing global crystal characteristics compared to previous successes of GNNs at capturing localized atomic environments 171. The CrystalTransformer model embodies this advance by feeding in crystal structure information through a Transformer encoder; specifically, this work starts with a concatenation of chemical and structural feature matrices, and then processes the combined representation into a Transformer encoder. In the encoder, the multi-head self-attention identifies complex, long-range correlations between all atoms in a manner that is agnostic to their spatial proximity, enabling the model to learn highly context-aware atomic embeddings 170. Other variants, such as MatInFormer, have generalized this paradigm by learning the ”grammar of crystallography” through the tokenization of space group information and subsequently demonstrating the adaptability of this architecture for property prediction on a range of datasets 219. Increasingly, a key factor is the strategic use of linguistic input. For example, if natural language descriptions of a crystalline solid are used, then high-accuracy predictions of properties such as bandgap can be achieved, which seems not possible for some GNNs due to an inability to capture certain nuanced, global structural details such as space group symmetry and Wyckoff positions 306. The success of Transformer-based architectures in the processing of both symbolic structural information and text suggests that they operate as superior contextual encoders rather than purely topological encoders 352. This allows them to implicitly integrate learned chemical and structural knowledge, avoiding the limitations of localized message-passing schemes.
@@
The prediction of mechanical properties such as bulk modulus and shear modulus is important in materials design with targeted strength, durability, and resilience 172, 179. Various machine learning models have been widely adopted for predicting such mechanical properties, greatly reducing their dependence on simulation methods that are computationally intensive 207. More recently, domain-specific LLMs for materials science can be specialized in the predictions of complex mechanical properties, often by leveraging text-derived representations or hybrid text+structured inputs. For example, ElaTBot is an LLM for the prediction of the full elastic constant tensor from text inputs 224. Hence, the elastic tensor can be predicted simultaneously with the determination of properties such as bulk modulus at any finite temperature. Moreover, by incorporating general-purpose LLMs like GPT-4o and using the Retrieval-Augmented Generation technique, its prediction accuracy can be boosted, indicating again the power of combining domain-specific knowledge with the broad reasoning capabilities of generalist models when retrieval sources are curated and verifiable 226. This significantly accelerates the prediction process but, more importantly, can support early-stage inverse design by proposing candidate compositions with targeted elastic properties—subject to the fidelity of the predictive surrogate and downstream verification.
@@
Among the most basic and important thermodynamic properties is the formation energy, which dictates stability for a crystalline solid 63. Its prediction is thus of paramount importance in order for high-throughput screening to be performed on new, synthesizable materials. Machine learning models have thus been employed to directly predict it from crystal structures, with the key ingredient being sophisticated feature engineering to account for periodicity in solids 107. Subsequent deep learning models demonstrated the feasibility of material property inference from merely elemental composition and symmetry classification 152.
A particularly noticeable development in this domain has been the development of structure-agnostic models, which predict properties from stoichiometry alone. Using a message-passing architecture on a dense weighted graph representation of the chemical formula, the Roost framework learns material descriptors in an explicit-structure-free way 132. This is especially important in the early stage of materials discovery when little to no crystal structural information is known. Self-supervised pre-training methods have been shown to boost the performances of such models even further, with substantial improvements in the low-data regime 91. Furthermore, LLMs have been demonstrated to handle simple chemical formulas and predict complex chemical properties (e.g., formation energy), significantly reducing the computational cost of materials screening.
@@
Thermoelectric materials enable the conversion between thermal and electrical energy, making them crucial for waste heat recovery and solid-state cooling technologies. However, the lack of large, structured datasets with relevant performance metrics hinders the development of models for high-performance thermoelectric materials 380. LLM-based extraction approaches are increasingly used to mitigate this data bottleneck. Recent work has shown that agentic, LLM-driven workflows autonomously extract the thermoelectric and structural properties from thousands of scientific articles 127. These intelligent agents can parse text, tables, and captions to create large machine-readable datasets that couple performance metrics with structural information such as the crystal class and space group 193. This automated curation pipeline can improve downstream regression by expanding training data and improving schema consistency; however, it still primarily optimizes proxy prediction objectives unless coupled to closed-loop validation and decision-making in the discovery loop.